# Cupid: Pose-Grounded Generative 3D Reconstruction from a Single Image

### v1

@@ Abstract
This work proposes a new generation-based 3D reconstruction method, named Cupid, that accurately infers the camera pose, 3D shape, and texture of an object from a single 2D image. Cupid casts 3D reconstruction as a conditional sampling process from a learned distribution of 3D objects, and it jointly generates voxels and pixel-voxel correspondences, enabling robust pose and shape estimation under a unified generative framework. By representing both input camera poses and 3D shape as a distribution in a shared 3D latent space, Cupid adopts a two-stage flow matching pipeline: (1) a coarse stage that produces initial 3D geometry with associated 2D projections for pose recovery; and (2) a refinement stage that integrates pose-aligned image features to enhance structural fidelity and appearance details. Extensive experiments demonstrate Cupid outperforms leading 3D reconstruction methods with an over 3 dB PSNR gain and an over 10% Chamfer Distance reduction, while matching monocular estimators on pose accuracy and delivering superior visual fidelity over baseline 3D generative models. For an immersive view of the 3D results generated by Cupid, please visit cupid3d.github.io.
Figure 1: Results for pose-grounded generative 3D reconstruction from a single test image. Given an input image (top left), Cupid estimates camera pose (bottom left) and reconstructs 3D model (bottom right), re-rendering the input (top right). It is robust to changes in scale, placement, and lighting while preserving fine details, and supports component-aligned scene reconstruction (bottom row). All results are produced in seconds via feed-forward sampling of the learned model. See cupid3d.github.io for an immersive view of the interactive 3D results.
@@ 1 Introduction
When we look at an object, we instinctively recognize the viewpoint from which it is seen—whether from the front, side, or back. This ability suggests that, based on a single view, humans can “reconstruct” a canonical 3D model of the object, including its shape and texture, and also estimate this viewpoint relative to the model (Sekuler & Palmer, 1992). Accurately determining both a canonical model of the object and an object-centric pose of the viewpoint is essential for effective interaction with the object (Gibson, 2014). For example, when we reach to grasp a cup, we can easily adjust the position of our hand based on our estimated viewpoint in relation to the cup’s canonical orientation. Therefore, it is crucial for embodied AI that robots develop similar abilities: from any given view, they should be able to (i) recover the 3D model of an object in its canonical pose (often referred to as the “prototypical” exemplar, which has a privileged status) and (ii) estimate the current camera viewpoint relative to the canonical object frame [^1] To simplify the presentation, “object” refers to the object model in its canonical pose, and “camera pose” refers to the object-centric camera pose in the following sections.. We call the task of estimating from images both an object’s canonical 3D model and object-centric camera poses as pose-grounded 3D reconstruction. Figure 1 illustrates the basic ideas with exemplar results produced by our new method.
Most existing learning-based 3D reconstruction or generation works only focus on one aspect of pose-grounded 3D reconstruction: either estimating relative poses between multiple input images (Wang et al., 2024; Zhang et al., 2024a), or reconstructing a view-centric 3D model without recovering the canonical model (Schönberger & Frahm, 2016; Hong et al., 2023), or generating 3D models without estimating the object-centric camera pose of the input image(s) (Xiang et al., 2025). To the best of our knowledge, no work has adequately addressed the two related objectives in the pose-grounded 3D reconstruction task within a unified framework.
Specifically, most multi-view based 3D reconstruction methods focus on recovering a 3D model for the observed parts from image(s) by leveraging correlations between images and view-centric depths (Yang et al., 2024), point clouds (Wang et al., 2025a; c), or radiance fields (Hong et al., 2023). When the poses of the input images are not provided, however, separate PnP or deep learning-based methods are often adopted for relative pose estimation (Wang et al., 2023; Xu et al., 2024c; Wang et al., 2024; Chen et al., 2022). Furthermore, all these methods approach reconstruction as a 2D-to-3D inverse problem based on the provided image(s), failing to account for the uncertainty caused by self-occluded or unseen regions. As a result, they struggle to accurately recover complete 3D objects or scenes from limited views. Ultimately, none of these reconstruction-based methods produces a canonical 3D model or estimates the viewpoint of each image relative to the canonical object pose, as desired for embodied AI applications.
Conversely, generative methods can synthesize canonically posed 3D objects (Lu et al., 2025) from a single image, offering priors over textures (Zhang et al., 2024b), parts (Yang et al., 2025b) and actionable labels (Liu et al., 2024a) that are difficult to obtain from reconstruction alone. However, current image-conditioned 3D generators typically infer plausible variations rather than produce a faithfully reconstructed 3D model based on the given image(s). As a result, they can exhibit shape inconsistencies and color shifts when real inputs deviate from the synthetic training distribution (Xiang et al., 2025). Furthermore, because pose is not explicitly modeled, integrating the so-generated results into 3D reconstruction typically requires costly pose optimization, and alignment remains fragile when the generated 3D object deviates from the observation (Yao et al., 2025; Dong et al., 2025). Thus, single-image (or few-image based) object reconstruction from a learned generative 3D prior remains largely an open challenge.
In this work, we address this challenge by identifying pose grounding as the key gap between 3D generation and reconstruction: to explain a single image, a model must recover dense 3D–2D correspondences between the object and the input viewpoint. To marry the benefits of reconstruction and generation, we introduce Cupid (for “in-Cube Pixel Distribution”), a pose-grounded generative framework that jointly models 3D voxels and 2D pixels within a unified framework. Cupid adopts a sparse voxel representation and a two-stage training scheme inspired by  Xiang et al. (2025). First, a flow model jointly samples 3D points and their 2D pixel correspondences, enabling efficient pose recovery via PnP (Abdel-Aziz et al., 2015). Second, conditioned on the recovered pose, Cupid refines geometry and appearance based on reprojection consistency. Unlike image-conditioned 3D generators that rely on global image features, Cupid back-projects pose-aligned locally-attended pixel features into the voxel grid, thereby improving geometric alignment and appearance fidelity. In essence, our method conducts Bayesian inference for the 3D reconstruction by leveraging the learned 3D generative distribution conditioned on the pose and visual cues of the input image, hence significantly closes the gap between generation and reconstruction. As Figure 1 shows, this combination of generative priors and observations allows Cupid to produce high-fidelity 3D reconstructions that are both geometrically accurate and texturally faithful to the input image.
The primary contributions of this work can be summarized as below:
  • A unified generative framework, Cupid, that grounds 3D synthesis in a single 2D image. For the first time, we jointly model 3D object and dense 2D-to-3D correspondences, ensuring the generated 3D object is explicitly tied to the input view.
  • A novel pose-conditioned refinement strategy. Our two-stage approach decouples pose estimation from the generation of detailed geometry and texture, a mechanism critical for achieving high-fidelity reconstructions with strong geometric and textural consistency.
  • State-of-the-art results across both reconstruction and pose estimation. Cupid sets a new standard for single-image reconstruction, improving fidelity by over 3 dB PSNR against top 3D reconstruction models, while achieving geometric accuracy competitive with top point-map estimators.
@@ 2.1 Problem definition and method overview
We formulate generative reconstruction as estimating the joint posterior $p(\mathcal{O},\bm{\theta}\mid\mathbf{I}^{\textrm{cond}})$ under the observation constraint $\mathbf{I}^{\textrm{cond}}=\mathbf{\mathcal{P}}(\mathcal{O},\bm{\theta})$, where $\mathbf{I}^{\textrm{cond}}$ is the input image, $\mathcal{O}$ is the 3D object, $\bm{\theta}$ is the object-centric camera pose, and $\mathbf{\mathcal{P}}(\cdot,\cdot)$ projects the 3D object to the image.
We first use an encoder $\varphi$ to map the 3D object and camera pose to a volumetric latent feature $\mathbf{z}=\varphi(\mathcal{O},\bm{\theta})$, which can be decoded into 3D representations (e.g., Gaussian splats or meshes) by specific decoders (Shen et al., 2023; Kerbl et al., 2023). We use Rectified Flow (Lipman et al., 2022) for the latent $\mathbf{z}$ generation. The model interpolates between data $\mathbf{z}_{0}$ and noise $\bm{\epsilon}$ over time $t$ via $\mathbf{z}_{t}=(1-t)\mathbf{z}_{0}+t\bm{\epsilon}$. The reverse process follows a time-dependent velocity field $\mathbf{v}(\mathbf{z}_{t},\mathbf{I}^{\textrm{cond}},t)=\nabla_{t}\mathbf{z}_{t}$ that drives noisy samples toward the data distribution. We parameterize $\mathbf{v}$ with a neural network $\mathbf{v}_{\phi}$ and train it using the Conditional Flow Matching (CFM) objective:
@@
We will first describe the representation of object and camera pose in Section 2.2. To enable efficient 3D generation, we utilize a cascaded flow modeling approach for $\mathbf{v}_{\phi}$ (Xiang et al., 2025). In the first stage, we generate an occupancy cube along with the camera pose. The second stage predicts the 3D shape and texture features in the occupied regions based on the outputs from the first stage. This process will be detailed in Section 2.3. The pipeline is illustrated in Figure 2.
Figure 2: Overview of Cupid. From a single input image, Cupid first generates an occupancy and a UV cube in canonical space. A Perspective-n-Point (PnP) solver (i.e., Equation 2) then recovers the camera pose. Using this recovered camera pose, we extract pose-aligned conditioning latents and visual features (i.e., Equation 3), along with noisy structured latents, to generate the geometry and appearance features, which will be decoded to the 3D Gaussian splats and mesh.
@@ 2.2 Representation of 3D Object and camera pose
To enable network training, we first tensorize 3D objects with a voxel-based representation: $\mathcal{O}\triangleq\{\mathbf{x}_{i},\mathbf{f}_{i}\}_{i=1}^{L}$, where $\mathbf{x}_{i}\in\mathbb{R}^{3}$ represents the coordinates of the $i$-th active voxel in a cube, and $\mathbf{f}_{i}$ encodes associated features derived from multi-view DINOv2 (Oquab et al., 2023) feature aggregation, subsequently compressed by a 3D variational autoencoder encoder (VAE). Here, $L$ denotes the number of voxels. This voxel set can be decoded back to 3D formats such as mesh or Gaussian splat (Xiang et al., 2025).
@@
Camera pose $\bm{\theta}$ is represented by the projection matrix $\mathbf{P}=\mathbf{K}[\mathbf{R}|\mathbf{t}]\in\mathbb{R}^{3\times 4}$, where $\mathbf{K}$ is intrinsic, and $(\mathbf{R},\mathbf{t})$ are extrinsic parameters mapping object to camera space. Inspired by recent pose estimation work (Wang et al., 2024), we propose a dense 3D pose representation [^2] While one can use a 12D vector for pose, such low-dimension parameterization can hinder neural optimization (Zhang et al., 2024a). via an in-cube pixel distribution (Cupid). Specifically, we reparameterize $\bm{\theta}$ as dense 3D-2D correspondence: $\bm{\theta}\triangleq\{\mathbf{x}_{i},\mathbf{u}_{i}\}_{i=1}^{L}$ where $\mathbf{u}_{i}:(u_{i},v_{i})\in[0,1]^{2}$ is the normalized 2D pixel coordinates anchored at the 3D voxel $\mathbf{x}_{i}\in\mathcal{O}$. Each anchor’s pixel coordinate is obtained by $\mathbf{u}_{i}=\pi\big(\mathbf{P},\mathbf{x}_{i}\big)$, where $\pi$ denotes perspective projection. Such reparameterized pose, i.e., 3D-2D correspondence can be presented as 3D UV cube. Given these correspondences, we recover the global camera matrix $\mathbf{P}^{*}$ using a least-squares solver (Abdel-Aziz et al., 2015):
@@
We then decompose $\mathbf{P}^{*}$ into $(\mathbf{K},\mathbf{R},\mathbf{t})$ via RQ decomposition. Unlike image-based pose representations (e.g., 2D ray or point maps (Zhang et al., 2024a; Wang et al., 2024)), we reparameterize pose within a 3D cube. Intuitively, the $(u,v)$ coordinates act like view-dependent colors defined on a 3D occupancy grid (see the UV cube in the Figure 2). Under this view, joint object–pose generation transforms to producing a 3D object with view-dependent colors.

### v2

# Cupid: Generative 3D Reconstruction via Joint Object and Pose Modeling

@@ Abstract
We introduce Cupid, a generative 3D reconstruction framework that jointly models the full distribution over both canonical objects and camera poses. Our two-stage flow-based model first generates a coarse 3D structure and 2D-3D correspondences to estimate the camera pose robustly. Conditioned on this pose, a refinement stage injects pixel-aligned image features directly into the generative process, marrying the rich prior of a generative model with the geometric fidelity of reconstruction. This strategy achieves exceptional faithfulness, outperforming state-of-the-art reconstruction methods by over 3 dB PSNR and 10% in Chamfer Distance. As a unified generative model that decouples the object and camera pose, Cupid naturally extends to multi-view and scene-level reconstruction tasks without requiring post-hoc optimization or fine-tuning.
footnotetext: $\dagger$} Equal contribution. $\ddagger$} Corresponding author
@@ 1 Introduction
Inferring an object’s 3D structure from a single 2D image is a fundamental challenge in computer vision. Consider the physics of image formation: $\mathbf{I}=\mathbf{\mathcal{P}}(\mathcal{O},\bm{\theta})$, where $\mathbf{I}$ is the rendered image, $\mathcal{O}$ is the 3D object in an object-centric, view-agnostic canonical frame, $\bm{\theta}$ is the camera pose in that frame, and $\mathbf{\mathcal{P}}(\cdot,\cdot)$ projects the 3D object onto the image. Since object and pose jointly determine $\mathbf{I}$, given a single image, it is extremely challenging to disentangle the intrinsic properties of object $\mathcal{O}$ (e.g., shape and texture) from the extrinsic property of the partial observation, namely the camera viewpoint $\bm{\theta}$.
Current approaches often tackle this problem from only one aspect. On one hand, 3D generative models [xiang2025structured, zhang2024clay, zhao2025hunyuan3d, liu2024cage, lu2025orientation] excel at generating a high-quality canonical 3D object $\mathcal{O}$, while completely ignoring the camera pose $\bm{\theta}$, a crucial component of image formation. As a result, contemporary image-conditioned 3D generators typically produce plausible variations, but struggle to achieve a faithful reconstruction consistent with the input view, leading to shape inconsistencies and color drift, especially when real-world inputs deviate from the training distribution. Furthermore, integrating generated 3D assets into complex scene reconstruction is highly challenging, requiring costly and fragile post-hoc optimization for spatial alignment [yao2025castcomponentaligned3dscene, dong2025hiscene, geng2025one]. The difficulty arises primarily from the missing camera pose prior, which is never modeled in these approaches.
On the other hand, common 3D reconstruction methods [schoenberger2016sfm, depth_anything_v2, wang2025vggt, wang2024dust3r, wang2025moge] produce 3D geometry through 2D-to-3D lifting in a pixel-aligned, view-centric frame, which essentially fixes the camera pose $\bm{\theta}$ as an identity constant. As a result, they lack the generative ability to plausibly “imagine” occluded regions, yielding incomplete models or collapsing the scene into a single, undifferentiated mesh even given sufficient views [wang2025vggt, wang2024dust3r, wang2025moge]. Although some methods [hong2023lrm, wang2023pf] attempt to deterministically complete unseen regions, learning such an entangled data distribution limits their ability to truly “understand” the underlying 3D spatial relationship that is crucial to embodied AI.
Figure 1: Generative 3D reconstruction from a single image. Left: Input image. Middle: Our method recovers a canonical 3D model (intrinsic properties such as shape and texture) along with the exact camera pose (extrinsic property) that generated the input. Right: Re-rendering the 3D model from the estimated pose faithfully reproduces the input.
Figure 2: Results for generative 3D reconstruction from a single test image. Given an input image (top left), Cupid estimates camera pose (bottom left) and reconstructs 3D model (bottom right), re-rendering the input (top right). It is robust to changes in scale, placement, and lighting while preserving fine details, and supports component-aligned scene reconstruction (bottom row). All results are produced in seconds via feed-forward sampling of the learned model. See cupid3d.github.io for an immersive view of the interactive 3D results.
Observing the gap between these two lines of work, we argue an ideal system should jointly model both components within a unified framework, deriving this joint distribution from a large set of 2D observations without fixing camera pose as identity or specifying how the canonical object is oriented***3D artist-curated training data is typically placed in a display-ready frame that naturally encodes human priors on canonicalness. Our learned representation captures this dominant mode without explicit labeling. . This mirrors how humans perceive the world: as illustrated in Fig. 1, we readily infer that the image captures the plush toy from its left frontal side. This implies that, through exposure to abundant visual data, we maintain a view-agnostic mental impression of 3D objects while simultaneously recalling viewpoint that would most reproduce a given 2D observation. This formulation essentially bridges 3D reconstruction and 3D generation into a unified task, which we call Generative 3D Reconstruction.
To this end, we introduce Cupid (for “in-CUbe PIxel Distribution”), a probabilistic reconstruction framework designed to achieve this synthesis by simultaneously generating a full 3D object $\mathcal{O}$ and reasoning the exact camera pose $\bm{\theta}$ that can faithfully reproduce the 2D observation. Cupid formulates reconstruction as a conditional sampling process within a unified flow-based model that operates in two stages. First, it jointly generates a coarse 3D structure and a dense field of 3D-to-2D correspondences, enabling camera pose recovery through a simple PnP solver [abdel2015direct]. Second, conditioned on this recovered pose, a refinement stage injects pixel-aligned features from the input image directly into the generative process. This novel conditioning mechanism enforces consistency with the visual evidence, marrying the rich prior of a generative model with the geometric fidelity of reconstruction. As Fig. 2 shows, this simple yet effective approach produces geometrically complete 3D reconstructions that are texturally faithful to the input image.
The primary contributions of this work can be summarized as follows:
  • A unified generative reconstruction framework. Cupid bridges 3D generation and reconstruction by jointly modeling a canonical 3D object and camera pose. This unified design excels at single-view reconstruction, and can be applied to multi-view and component-aligned scene reconstruction without task-specific tuning or optimization.
  • A novel pose-conditioned refinement strategy. Our two-stage approach uses the recovered pose to inject pixel-local features, a mechanism critical for achieving high-fidelity reconstructions free from the shape and color inconsistencies typical of generative models.
  • Strong object-level reconstruction performance. Cupid leverages its generative capabilities to set a new standard in reconstruction fidelity, outperforming top holistic reconstruction models [hong2023lrm] by over 3 dB PSNR while matching the geometric accuracy of state-of-the-art monocular estimators [wang2025vggt, wang2025moge].
2 Related Work
3D reconstruction from many images.
Complete 3D reconstruction traditionally requires multiple views and Structure-from-Motion [schoenberger2016sfm, schoenberger2016mvs]. DUSt3R [wang2024dust3r] predicts pixel-aligned point maps for image pairs, enabling efficient recovery of view-centric camera poses and partial geometry; subsequent works improve efficiency and flexibility [wang2025vggt, wang2025moge2, Qianqian_2025_CVPR, zhang2025flarefeedforwardgeometryappearance, Yang_2025_Fast3R, karaev24cotracker3, depth_anything_v2, slam3r, must3r_cvpr25]. CUT3R processes images recurrently [Qianqian_2025_CVPR]; VGGT jointly predicts depth, point maps, and poses [wang2025vggt]; and MoGe introduces affine-invariant point maps for monocular geometry [wang2025moge]. These approaches lack generative capabilities and cannot plausibly infer occluded regions, resulting in partial geometry from limited input views. Even with abundant views, they produce a single fused scene without object separation. In contrast, our method models a joint distribution over 3D objects and camera pose, enabling completion of partial observations into fully compositional, object-level scenes.
3D reconstruction from one or a few images.
High-quality multiview data are hard to obtain in practice, motivating single- or sparse-view reconstruction with large models [xu2024sparp, wei2024meshlrm, liu2024meshformer, TripoSR2024, tang2024lgm, xu2024grm, xu2024freesplatter]. LRM [hong2023lrm] shows that complete 3D can be regressed from a single image under a fixed camera. PF-LRM [wang2023pf] enables pose-free sparse-view reconstruction via differentiable PnP [chen2022epro]. Subsequent work improves single-view quality by augmenting novel views from 2D generative priors and using more efficient representations [xu2024instantmesh, xu2024grm, liu2023one2345] or architectures [LaRa]. Yet their entangled pixel-aligned representations never decouple 3D content and viewpoint, typically assuming a fixed camera (e.g., unit focal length), making them poorly suited to real-world images with unknown camera parameters. Their deterministic formulations cannot model uncertainty in unseen regions, failing to produce plausible 3D variations. Our generative framework instead enables synthesizing diverse, high-fidelity 3D objects under flexible conditioning.
Figure 3: Overview of Cupid’s pipeline. From a single input image, Cupid first generates an occupancy and a UV cube in canonical space. Then, a Perspective-n-Point (PnP) solver (i.e., Eq. 2) recovers the camera pose. Using this recovered camera pose, we extract pose-aligned conditioning latents and visual features (i.e., Eq. 3), along with noisy structured latents, to generate the geometry and appearance features, which will be decoded to the 3D Gaussian splats and mesh.
3D reconstruction from generative priors.
Dreamfusion [poole2022dreamfusion] pioneers 3D generation from 2D diffusion models [rombach2021highresolution, saharia2022photorealistic]. Zero-1-to-3 [liu2023zero1to3] and subsequent work [shi2023zero123++, watson2022novel, li2024instant3d, liu2024syncdreamer, gao2024cat3d, hoellein2024viewdiff, wu2024reconfusion] fine-tune image diffusion models for single image-to-3D generation. Recent methods [zhou2025stable, ren2025gen3c, song2025historyguidedvideodiffusion] leverage video diffusion priors to achieve greater view consistency. Nevertheless, they rely on 2D generative priors that require dense supervision to reconstruct 3D objects and inherit the limitations of many-image-to-3D, most notably their difficulty in producing compositional scenes. Recent work [huang2025midi, yao2025castcomponentaligned3dscene, dong2025hiscene] incorporate native 3D generative priors [xiang2025structured, zhang2024clay] for scene reconstruction. Yet, without explicitly modeling camera pose, these methods resort to costly and fragile post-hoc optimization or additional tuning on spatial layouts [huang2025midi]. Notably, CAST [yao2025castcomponentaligned3dscene] designs a multi-stage pipeline for spatial alignment: it first estimates a point cloud [wang2025moge] then aligns each generated object via 3D–3D correspondences. In contrast, our approach achieves image-to-3D alignment in a single forward pass while delivering monocular geometry quality on par with MoGe [wang2025moge], enabling streamlined compositional scene reconstruction without complex pipelines.
@@ 3.1 Problem definition and method overview
We formulate our generative reconstruction as estimating the joint posterior $p(\mathcal{O},\bm{\theta}\mid\mathbf{I}^{\textrm{cond}})$ under the observation constraint $\mathbf{I}^{\textrm{cond}}=\mathbf{\mathcal{P}}(\mathcal{O},\bm{\theta})$, where $\mathbf{I}^{\textrm{cond}}$ is the input image, $\mathcal{O}$ is the 3D object, $\bm{\theta}$ is the object-centric camera pose, and $\mathbf{\mathcal{P}}(\cdot,\cdot)$ projects the 3D object to the image.
We first use an encoder $\varphi$ to map the 3D object and camera pose to a volumetric latent feature $\mathbf{z}=\varphi(\mathcal{O},\bm{\theta})$, which can be decoded into 3D representations (e.g., Gaussian splats or meshes) by specific decoders [shen2023flexible, kerbl20233d]. We use Rectified Flow [lipman2022flow] for the latent $\mathbf{z}$ generation. The model interpolates between data $\mathbf{z}_{0}$ and noise $\bm{\epsilon}$ over time $t$ via $\mathbf{z}_{t}=(1-t)\mathbf{z}_{0}+t\bm{\epsilon}$. The reverse process follows a time-dependent velocity field $\mathbf{v}(\mathbf{z}_{t},\mathbf{I}^{\textrm{cond}},t)=\nabla_{t}\mathbf{z}_{t}$ that drives noisy samples toward the data distribution. We parameterize $\mathbf{v}$ with a neural network $\mathbf{v}_{\phi}$ and train it using the Conditional Flow Matching (CFM) objective:
@@
We first describe the representation of objects and camera poses in Sec. 3.2. To enable efficient 3D generation, we utilize a cascaded flow modeling approach for $\mathbf{v}_{\phi}$ [xiang2025structured]. In the first stage, we generate an occupancy cube along with the camera pose. The second stage predicts the 3D shape and texture features in the occupied regions based on the outputs from the first stage. This process is detailed in Sec. 3.3. The pipeline is illustrated in Fig. 3.
@@ 3.2 Representation of 3D Object and camera pose
To enable network training, we first tensorize 3D objects with a voxel-based representation: $\mathcal{O}\triangleq\{\mathbf{x}_{i},\mathbf{f}_{i}\}_{i=1}^{L}$, where $\mathbf{x}_{i}\in\mathbb{R}^{3}$ represents the coordinates of the $i$-th active voxel in a cube, and $\mathbf{f}_{i}$ encodes associated features derived from multi-view DINOv2 [oquab2023dinov2] feature aggregation, subsequently compressed by a 3D variational autoencoder (VAE) encoder. Here, $L$ denotes the number of voxels. This voxel set can be decoded to 3D formats such as mesh or Gaussian splat [xiang2025structured].
@@
Camera pose $\bm{\theta}$ is represented by the projection matrix $\mathbf{P}=\mathbf{K}[\mathbf{R}|\mathbf{t}]\in\mathbb{R}^{3\times 4}$, where $\mathbf{K}$ is intrinsic, and $(\mathbf{R},\mathbf{t})$ are extrinsic parameters mapping object to camera space. Inspired by recent pose estimation work [wang2024dust3r, zhang2024cameras], we propose an over-parameterized pose representationWhile camera pose can be encoded as a compact 1D token (e.g., a 12D vector), it can be challenging for 3D generators with 3D tokens [zhang2024cameras]. via an in-cube pixel distribution. Specifically, we reparameterize $\bm{\theta}$ as dense 3D-2D correspondence: $\bm{\theta}\triangleq\{\mathbf{x}_{i},\mathbf{u}_{i}\}_{i=1}^{L}$ where $\mathbf{u}_{i}:(u_{i},v_{i})\in[0,1]^{2}$ is the normalized 2D pixel coordinates anchored at the 3D voxel $\mathbf{x}_{i}\in\mathcal{O}$. Each anchor’s pixel coordinate is obtained by $\mathbf{u}_{i}=\pi\big(\mathbf{P},\mathbf{x}_{i}\big)$, where $\pi$ denotes perspective projection. Such a reparameterized pose, i.e., 3D-2D correspondence, can be presented as a 3D UV cube. Given these correspondences, we recover the global camera matrix $\mathbf{P}^{*}$ using a least-squares solver [abdel2015direct]:
@@
We then decompose $\mathbf{P}^{*}$ into $(\mathbf{K},\mathbf{R},\mathbf{t})$ via RQ decomposition. Unlike image-based pose representations (e.g., 2D ray or point maps [zhang2024cameras, wang2024dust3r]), we reparameterize pose within a 3D cube. Intuitively, the $(u,v)$ coordinates act like view-dependent colors defined on a 3D occupancy grid (see the UV cube in the Fig. 3). Under this view, joint object–pose generation transforms to producing a 3D object with view-dependent colors.